# Less is More: Improving LLM Alignment via Preference Data Selection

### v3

@@ 1 Introduction
Reinforcement Learning from Human Feedback (RLHF; christiano2017deep, ; ziegler2019fine, ) has emerged as a crucial technique for aligning Large Language Models (LLMs) with human preferences and values. Traditional RLHF implementations involve a two-stage process: reward model training based on preference data followed by reinforcement learning optimization. However, this approach presents significant computational challenges, requiring loading multiple model instances and extensive hyperparameter tuning.
As an alternative, (rafailov2024direct, ) introduced Direct Preference Optimization (DPO), which streamlines the alignment process by directly optimizing the LLM policy from preference data. DPO has demonstrated comparable effectiveness while substantially reducing computational requirements compared to classical RLHF. Following DPO’s introduction, numerous studies have proposed improvements through modified learning objectives (zhao2023slic, ; azar2024general, ; ethayarajh2024kto, ) and iterative learning schemes (xiong2024iterative, ). While these algorithmic advances have shown promise, there remains a critical gap in our understanding of the data-centric aspects of preference learning: what characteristics of preference data contribute most to model alignment?
@@ 1.1 Related Work
Preference learning algorithms. Reinforcement Learning from Human Feedback also known as dueling RL (pacchiano2021dueling, ) or preference-based RL (chen2022human, ), has become a crucial component of recent Large Language Models (LLMs) such as ChatGPT (ouyang2022training, ). While the classical RLHF pipeline traditionally uses Proximal Policy Optimization, several alternative approaches have been proposed. These include but not limited other RL-based training algorithms (li2023remax, ; zhong2024dpo, ), rejection sampling (dong2023raft, ; gulcehre2023reinforced, ), conditional supervised fine-tuning (lu2022quark, ; yang2024rewards, ; zhang2024reward, ), and Direct Preference Optimization (rafailov2024direct, ). Among these alternatives, DPO has gained significant attention due to its simplicity and robust performance. Following the introduction of DPO, numerous works (zhao2023slic, ; azar2024general, ; ethayarajh2024kto, ; meng2024simpo, ; tang2024generalized, ; han2024f, ; xu2024contrastive, ; hong2024orpo, ; wu2024self, ) have attempted to improve its performance by modifying the DPO objective.
Data selection in LLM Fine-tuning. Data selection is crucial in LLM post-training (wang2024survey, ) for two key observations: post-training typically converges rapidly, and excessive data can degrade model performance through overfitting or exposure to toxic content (swayamdipta2020dataset, ; deng2023counterfactual, ). Recent research has focused on enhancing instruction tuning efficiency by identifying high-quality subsets from large instruction datasets (cao2023instruction, ; li2024superfiltering, ; xia2024less, ), often adapting active learning query strategies (ren2021survey, ) to assess sample uncertainty and diversity. However, data efficiency in preference learning remains relatively unexplored. Prior studies have studied reducing annotation costs in preference dataset creation (muldrew2024active, ; yasunaga2024alma, ) and on scenarios involving numerous ranking annotations (thekumparampil2024comparing, ; mukherjee2024optimal, ). Other research aims to improve a model’s ability to distinguish between two responses by adding a margin to the loss term (meng2024simpo, ; park2024disentangling, ; azar2024general, ). Additionally, concurrent work highlights the importance of margins for preference data filtering, though there is debate on whether hard samples help or hinder preference learning (wu2024beta, ; khaki2024rs, ; yu2025rip, ; gao2025principled, ).
Our work firstly provides clear criteria for identifying informative samples while filtering toxic ones, thereby improving both DPO’s efficiency and performance. Furthermore, our method extends to iterative DPO (xiong2024iterative, ) and its variants (zhang2024self, ), wherein training data is dynamically generated by the model during its iterative training process.
@@ 2 Background
Reward Learning. In the reward learning stage, a reward model is trained to approximate human preferences based on preference data. By adopting the Bradley-Terry model (bradley1952rank, ) to capture human preference, reward training involves minimizing the loss:
@@
Policy Optimization with Reinforcement Learning. Once the reward model $r$ is trained, it is used to guide the optimization of a policy $\pi_{\theta}(y|x)$, where $\theta$ denotes the parameters of the model. This stage often employs reinforcement learning techniques such as Proximal Policy Optimization (PPO; schulman2017proximal, ) to optimize the policy by maximizing the expected reward.
@@
Recently, as an alternative to the RL-based policy optimization in RLHF, Direct Preference Optimization (DPO; rafailov2024direct, ) has been proposed. DPO simplifies the reward alignment process by directly incorporating human preference data into supervised training. Instead of defining and optimizing a reward function explicitly, DPO minimizes
@@ 3.1 Parameter Shrinkage and Inflation Analysis
We follow the model from (zhu2023principled, ) to illustrate why data selection can improve model performance. We assume that reward model $r(x,y)=\langle\phi(x,y),\omega^{*}\rangle$ with some feature function $\phi(\cdot,\cdot)$. For reward learning, our reward model can be an explicit $r(x,y)$ (ouyang2022training, ), while for DPO, $\beta\log\frac{\pi_{\theta}(y|x)}{\pi_{\mathrm{ref}}(y|x)}$ plays the role of reward model implicitly (rafailov2024direct, ). Based on observations in previous literature, we can derive such features by removing the last layer of the pre-trained model. However, both humans and other large models may use inaccurate reward functions to generate labels, where the labels represent the ranking of two responses. We say the preference between $y_{w}$ and $y_{l}$ is generated by $r(x,y_{w})-r(x,y_{l})+\zeta$ where $\zeta$ is an exogenous error. We use $\Delta\phi(x)$ to denote $\phi(x,y_{w})-\phi(x,y_{l})$ for simplicity.
@@
Since the optimal $\omega$ is $\omega^{*}$ without $\zeta$, plugging $\omega^{*}$ in Equation 2 will cause the left-hand side to be greater than the right-hand side. Therefore, the optimal $\omega$ with the existence of $\zeta$ intends to shrink to the original point compared to $\omega^{*}$ so that the first-order condition is still satisfied.
We provide the underlying intuition with an extreme example. If $\mathbb{V}(\zeta)$ goes to infinity, the preference between $y_{w}$ and $y_{l}$ mainly depends on $\zeta$, approaching a Rademacher distribution, then $\omega=0$ could be a good solution to Equation 1. In other words, $\zeta$ offsets part of the information provided by the reward model, causing the model’s parameters to shrink toward zero. Thus, data selection is essential for acquiring policies with good performance. Finally, we remark that $\zeta$ can come from multiple resources, including human classification errors, different embeddings or reward models from other LLMs and so on.
@@
First, when the margin is large, namely, $\langle\Delta\phi(x),\omega^{*}\rangle+\zeta$ is large, from the S-shaped graph of $\sigma(\cdot)$, we know that the slope is very small in this area. As a result, the probability of preference reversal caused by $\zeta$ is low, which means the likelihood of incorrect samples is also low. Secondly, given prompt $x$, as we select data with large $\langle\Delta\phi(x),\omega^{*}\rangle+\zeta$, the posterior distribution of $\zeta$ is skewed toward the positive side. Therefore, the preferences corresponding to this kind of data are more pronounced, leading to inflated estimates of $\omega$ in Equation 1. Finally, we point out that if realized $y_{w}$ and $y_{l}$ are all separable, proportional scaling of $\omega$ can reduce the value of Equation 1 continuously. Hence, some techniques like regularization or early stopping when training are indispensable.
@@ 3.2 Multi-source Margin Aggregation
Building on our previous analysis, we aim to develop a data selection strategy based on the margin-maximization principle, with the calculation of reward margin being the critical component. We examine two distinct types of reward margin calculations: external reward margin and implicit reward margin. The external reward margin is provided by an external reward function, while the implicit reward margin is derived from the implicit reward $\log\frac{\pi_{\theta}}{\pi_{\mathrm{ref}}}$, where $\pi_{\theta}$ represents the policy trained by DPO (rafailov2024direct, ) (see Section 4.1 for details on margin calculation). We visualize the joint distribution of different reward margin sources using the UltraFeedback dataset (cui2023ultrafeedback, ) in Figure LABEL:fig:visual.
  • The left and middle panels of Figure LABEL:fig:visual reveal several key phenomena: (1) The correlation between implicit and external reward margins is notably weak. In particular, samples exhibiting high positive implicit margins span a broad range of external margins, including strongly negative values, and vice versa. This divergence highlights the distinct preference patterns captured by these two reward types. This underscores the need to combine both reward types for a reliable margin estimation. (2) In contrast, we observe a strong correlation between implicit reward margins calculated by models of different sizes (Llama-3.2 3B and 1B). Notably, these two patterns remain consistent across other datasets as well (see Appendix B.2).
  • Online RLHF (xiong2024iterative, ; dong2025self, ) employs the target model to generate multiple responses for given prompts iteratively and uses an external reward model to identify the response pair with the largest margin for DPO training. The right panel of Figure LABEL:fig:visual illustrates that a max-margin pair construction method, even when derived from one strong reward model, can still yield ambiguous preferences when evaluated by another reward model (which shows similar performance on RewardBench Leadboard). This ambiguity in preference signal indicates that the online data generation process can still cause high redundancy, which may offer little to no benefit, or could even be detrimental, to online-DPO training.
@@
Step 1: in-distribution pre-DPO training. Our objective is to obtain the in-distribution implicit reward model with low computational and GPU-memory cost. Given the strong correlation of that margin across different models (See Figure LABEL:fig:visual), this weak-to-strong guidance is feasible. To achieve this, we randomly select a small seed dataset $\mathcal{D}_{0}$ from $\mathcal{D}$ and employ a (or several) small model to perform preference learning on this seed set. The high sample and training efficiency of the DPO loss (kim2025spread, ) ensures the feasibility of this approach.
@@
Step 3: Bayesian aggregation. To mitigate noise from individual reward sources, we utilize Bayesian probability theory for their robust aggregation. Given that reward margins from different sources often vary in their underlying distributions and value ranges, we propose projecting these diverse margins into a unified probability space. This projection serves to quantify the confidence that a specific preference direction $y_{w}>y_{l}$ is correct, formally expressed as $\mathbb{P}(y_{w}>y_{l}|m^{1},m^{2},\cdots,m^{K})$. Assuming that these sources are conditionally independent, then following the Bayesian formula transformation in previous work (liu2022mpc, ; deng2024a3s, ), the preference probability can be expressed as:
@@ 4 Experiments
Table 1: GPT-4 judged win rates for Llama-3.2-3B models fine-tuned with DPO on subsets selected by various data selection strategies. For every strategy and benchmark (TL;DR, HH, UltraFeedback (UF)), 2,000 preference samples were selected. Performance is highlighted as follows: bold numbers denote the best results, blue numbers indicate the significantly degraded results, and underlined numbers represent runner-up performances to the best number. P, Z, and N denote the most positive, near-zero, and most negative selection principles.

### v4

@@ 1 Introduction
Reinforcement Learning from Human Feedback (RLHF; Christiano et al., 2017; Ziegler et al., 2019) has emerged as a crucial technique for aligning Large Language Models (LLMs) with human preferences and values. Traditional RLHF implementations involve a two-stage process: reward model training based on preference data followed by reinforcement learning optimization. However, this approach presents significant computational challenges, requiring loading multiple model instances and extensive hyperparameter tuning.
As an alternative, (Rafailov et al., 2024) introduced Direct Preference Optimization (DPO), which streamlines the alignment process by directly optimizing the LLM policy from preference data. DPO has demonstrated comparable effectiveness while substantially reducing computational requirements compared to classical RLHF. Following DPO’s introduction, numerous studies have proposed improvements through modified learning objectives (Zhao et al., 2023; Azar et al., 2024; Ethayarajh et al., 2024) and iterative learning schemes (Xiong et al., 2024). While these algorithmic advances have shown promise, there remains a critical gap in our understanding of the data-centric aspects of preference learning: what characteristics of preference data contribute most to model alignment?
@@ 1.1 Related Work
Preference learning algorithms. Reinforcement Learning from Human Feedback also known as dueling RL (Pacchiano et al., 2021) or preference-based RL (Chen et al., 2022), has become a crucial component of recent Large Language Models (LLMs) such as ChatGPT (Ouyang et al., 2022). While the classical RLHF pipeline traditionally uses Proximal Policy Optimization, several alternative approaches have been proposed. These include but not limited other RL-based training algorithms (Li et al., 2023c; Zhong et al., 2024), rejection sampling (Dong et al., 2023; Gulcehre et al., 2023), conditional supervised fine-tuning (Lu et al., 2022; Yang et al., 2024b; Zhang et al., 2024a), and Direct Preference Optimization (Rafailov et al., 2024). Among these alternatives, DPO has gained significant attention due to its simplicity and robust performance. Following the introduction of DPO, numerous works (Zhao et al., 2023; Azar et al., 2024; Ethayarajh et al., 2024; Meng et al., 2024; Tang et al., 2024; Han et al., 2024; Xu et al., 2024; Hong et al., 2024; Wu et al., 2024c) have attempted to improve its performance by modifying the DPO objective.
Data selection in LLM Fine-tuning. Data selection is crucial in LLM post-training (Wang et al., 2024) for two key observations: post-training typically converges rapidly, and excessive data can degrade model performance through overfitting or exposure to toxic content (Swayamdipta et al., 2020; Deng et al., 2023b). Recent research has focused on enhancing instruction tuning efficiency by identifying high-quality subsets from large instruction datasets (Cao et al., 2024; Li et al., 2024; Xia et al., 2024), often adapting active learning query strategies (Ren et al., 2021) to assess sample uncertainty and diversity. However, data efficiency in preference learning remains relatively unexplored. Prior studies have studied reducing annotation costs in preference dataset creation (Muldrew et al., 2024; Yasunaga et al., 2024) and on scenarios involving numerous ranking annotations (Thekumparampil et al., 2024; Mukherjee et al., 2024). Other research aims to improve a model’s ability to distinguish between two responses by adding a margin to the loss term (Meng et al., 2024; Park et al., 2024; Azar et al., 2024). Additionally, concurrent work highlights the importance of margins for preference data filtering, though there is debate on whether hard samples help or hinder preference learning (Wu et al., 2024b; Khaki et al., 2024; Yu et al., 2025; Gao et al., 2025).
Our work firstly provides clear criteria for identifying informative samples while filtering toxic ones, thereby improving both DPO’s efficiency and performance. Furthermore, our method extends to iterative DPO (Xiong et al., 2024) and its variants (Zhang et al., 2024b), wherein training data is dynamically generated by the model during its iterative training process.
@@ 2 Background
Reward Learning. In the reward learning stage, a reward model is trained to approximate human preferences based on preference data. By adopting the Bradley-Terry model (Bradley and Terry, 1952) to capture human preference, reward training involves minimizing the loss:
@@
Policy Optimization with Reinforcement Learning. Once the reward model $r$ is trained, it is used to guide the optimization of a policy $\pi_{\theta}(y|x)$, where $\theta$ denotes the parameters of the model. This stage often employs reinforcement learning techniques such as Proximal Policy Optimization (PPO; Schulman et al., 2017) to optimize the policy by maximizing the expected reward.
@@
Recently, as an alternative to the RL-based policy optimization in RLHF, Direct Preference Optimization (DPO; Rafailov et al., 2024) has been proposed. DPO simplifies the reward alignment process by directly incorporating human preference data into supervised training. Instead of defining and optimizing a reward function explicitly, DPO minimizes
@@ 3.1 Parameter Shrinkage and Inflation Analysis
Figure 1: The workflow of the BeeS method.
We follow the model from (Zhu et al., 2023) to illustrate why data selection can improve model performance. We assume that reward model $r(x,y)=\langle\phi(x,y),\omega^{*}\rangle$ with some feature function $\phi(\cdot,\cdot)$. For reward learning, our reward model can be an explicit $r(x,y)$ (Ouyang et al., 2022), while for DPO, $\beta\log\frac{\pi_{\theta}(y|x)}{\pi_{\mathrm{ref}}(y|x)}$ plays the role of reward model implicitly (Rafailov et al., 2024). Based on observations in previous literature, we can derive such features by removing the last layer of the pre-trained model. However, both humans and other large models may use inaccurate reward functions to generate labels, where the labels represent the ranking of two responses. We say the preference between $y_{w}$ and $y_{l}$ is generated by $r(x,y_{w})-r(x,y_{l})+\zeta$ where $\zeta$ is an exogenous error. We use $\Delta\phi(x)$ to denote $\phi(x,y_{w})-\phi(x,y_{l})$ for simplicity.
@@
Since the optimal $\omega$ is $\omega^{*}$ without $\zeta$, plugging $\omega^{*}$ in Equation˜2 will cause the left-hand side to be greater than the right-hand side. Therefore, the optimal $\omega$ with the existence of $\zeta$ intends to shrink to the original point compared to $\omega^{*}$ so that the first-order condition is still satisfied.
We provide the underlying intuition with an extreme example. If $\mathbb{V}(\zeta)$ goes to infinity, the preference between $y_{w}$ and $y_{l}$ mainly depends on $\zeta$, approaching a Rademacher distribution, then $\omega=0$ could be a good solution to Equation˜1. In other words, $\zeta$ offsets part of the information provided by the reward model, causing the model’s parameters to shrink toward zero. Thus, data selection is essential for acquiring policies with good performance. Finally, we remark that $\zeta$ can come from multiple resources, including human classification errors, different embeddings or reward models from other LLMs and so on.
Table 1: Symbols used in the formulation.
r x $y_{w}/y_{l}$ $\phi/\Delta\phi$ $\zeta$ w
Reward Input prompt Preferred/dispreferred Response (relative) feature function Exogenous error Learnable parameters
@@
First, when the margin is large, namely, $\langle\Delta\phi(x),\omega^{*}\rangle+\zeta$ is large, from the S-shaped graph of $\sigma(\cdot)$, we know that the slope is very small in this area. As a result, the probability of preference reversal caused by $\zeta$ is low, which means the likelihood of incorrect samples is also low. Secondly, given prompt $x$, as we select data with large $\langle\Delta\phi(x),\omega^{*}\rangle+\zeta$, the posterior distribution of $\zeta$ is skewed toward the positive side. Therefore, the preferences corresponding to this kind of data are more pronounced, leading to inflated estimates of $\omega$ in Equation˜1. Finally, we point out that if realized $y_{w}$ and $y_{l}$ are all separable, proportional scaling of $\omega$ can reduce the value of Equation˜1 continuously. Hence, some techniques like regularization or early stopping when training are indispensable.
@@ 3.2 Multi-source Margin Aggregation
(a) Visualization of joint margin distribution on UltraFeedback. (Left) Joint distribution of external and implicit reward margin values. (Middle) Joint distribution of implicit reward margins computed using models of 1B and 3B scales. (Right) Joint distribution of two different external reward margin values on online-generated data.
Building on our previous analysis, we aim to develop a data selection strategy based on the margin-maximization principle, with the calculation of reward margin being the critical component. We examine two distinct types of reward margin calculations: external reward margin and implicit reward margin. The external reward margin is provided by an external reward function, while the implicit reward margin is derived from the implicit reward $\log\frac{\pi_{\theta}}{\pi_{\mathrm{ref}}}$, where $\pi_{\theta}$ represents the policy trained by DPO (Rafailov et al., 2024) (see Section 4.1 for details on margin calculation). We visualize the joint distribution of different reward margin sources using the UltraFeedback dataset (Cui et al., 2024) in Figure 4(a).
  • The left and middle panels of Figure 4(a) reveal several key phenomena: (1) The correlation between implicit and external reward margins is notably weak. In particular, samples exhibiting high positive implicit margins span a broad range of external margins, including strongly negative values, and vice versa. This divergence highlights the distinct preference patterns captured by these two reward types. This underscores the need to combine both reward types for a reliable margin estimation. (2) In contrast, we observe a strong correlation between implicit reward margins calculated by models of different sizes (Llama-3.2 3B and 1B). Notably, these two patterns remain consistent across other datasets as well (see Appendix B.2).
  • Online RLHF (Xiong et al., 2024; Dong et al., 2025) employs the target model to generate multiple responses for given prompts iteratively and uses an external reward model to identify the response pair with the largest margin for DPO training. The right panel of Figure 4(a) illustrates that a max-margin pair construction method, even when derived from one strong reward model, can still yield ambiguous preferences when evaluated by another reward model (which shows similar performance on RewardBench Leadboard). This ambiguity in preference signal indicates that the online data generation process can still cause high redundancy, which may offer little to no benefit, or could even be detrimental, to online-DPO training.
@@
Step 1: in-distribution pre-DPO training. Our objective is to obtain the in-distribution implicit reward model with low computational and GPU-memory cost. Given the strong correlation of that margin across different models (See Figure 4(a)), this weak-to-strong guidance is feasible. To achieve this, we randomly select a small seed dataset $\mathcal{D}_{0}$ from $\mathcal{D}$ and employ a (or several) small model to perform preference learning on this seed set. The high sample and training efficiency of the DPO loss (Kim et al., 2025) ensures the feasibility of this approach.
@@
Step 3: Bayesian aggregation. To mitigate noise from individual reward sources, we utilize Bayesian probability theory for their robust aggregation. Given that reward margins from different sources often vary in their underlying distributions and value ranges, we propose projecting these diverse margins into a unified probability space. This projection serves to quantify the confidence that a specific preference direction $y_{w}>y_{l}$ is correct, formally expressed as $\mathbb{P}(y_{w}>y_{l}|m^{1},m^{2},\cdots,m^{K})$. Assuming that these sources are conditionally independent, then following the Bayesian formula transformation in previous work (Liu et al., 2022; Deng et al., 2024), the preference probability can be expressed as:
@@ 4 Experiments
4.1 Experimental Setup
Preference Datasets.
We evaluate our approach using three established preference datasets: (1) Reddit TL;DR summarization dataset (Völske et al., 2017; Stiennon et al., 2020) that contains human-written summaries and human-rated results, (2) Anthropic Helpful and Harmless dialogue dataset (HH) (Bai et al., 2022), and (3) UltraFeedback (Cui et al., 2024), which comprises quality-scored model responses across diverse prompts from multiple sources. To explore how models react to on-policy data, we leverage two modified versions of the UltraFeedback dataset, Llama-UltraFeedback and Mistral-UltraFeedback (Meng et al., 2024). In the variants, the original chosen and rejected responses are replaced with the highest and lowest scored responses, respectively, sampled from five candidates generated by the corresponding Instruct model. The scores are given by the PairRM (Jiang et al., 2023b) reward model. Statistics about these datasets are in Appendix A.
Models.
Our experiments are conducted across four model series: Llama-3.2 (MetaAI, 2024), Llama-3 (Dubey et al., 2024), Mistral-7B-v2 (Jiang et al., 2023a), and Qwen-2.5 (Yang et al., 2024a) under Base and Instruct setups. For the Base model (Llama-3.2-3B and Llama-3-8B), we first establish fundamental Instruction-following capabilities through supervised fine-tuning on the RLHFlow/SFT-OpenHermes-2.5-Standard datasets. For the Instruct setup, we directly use them as the start of DPO training. Regarding the external reward model, we adopt the recent Skywork-Reward-Llama-3.1-8B-v0.2 (Liu et al., 2024) that is the best reward model at this scale according to the RewardBench leadboard. As for the implicit reward model, we employ the Llama-3.2-3B Base and its DPO-tuned model (on 2,000 randomly selected samples from the complete dataset) for $\pi_{\mathrm{ref}}$ and $\pi_{\theta}$.
Implementation and Evaluation.
For DPO training, we follow (Rafailov et al., 2024) and use a fixed value of $\beta=0.1$, except for TL;DR where $\beta=0.5$. We run each training for two epochs, with a learning rate of $5\times 10^{-7}$, and a $0.1$ warmup ratio. Following (Rafailov et al., 2024), we evaluate the models using 400 randomly sampled test sets from the validation/test pools of the TL;DR and HH datasets, separately. For models trained on UltraFeedback, we employ AlpacaEval and AlpacaEval 2.0 (Li et al., 2023b) as our evaluation benchmark, which consists of 805 diverse questions. [^2] They use the same set of questions, and differ in their reference response generation: AlpacaEval uses Text-Davinci-003 (Ye et al., 2023), whereas AlpacaEval 2.0 employs GPT4-1106-preview As the ground truth oracle is unavailable, we employ GPT-4 as a proxy for human judgment across three distinct settings: summarization, helpful or harmless completion, and single-turn dialogue. We utilized a fixed decoding temperature ($T=0.7$) for all model generation in the experiments. More details are presented in Appendix A.
Baselines.
For the offline data selection setting, we compare our method with three types of methods: (1) Random, a simple yet effective strategy in many domains (e.g., Instruction tuning (Xia et al., 2024)), (2) IFD (Li et al., 2023a) (i.e., exponential form of the Point-wise Mutual Information), which measures semantic overlap. We use the difference in IFD scores between chosen and rejected responses for preference data selection. (3) External/Implicit Margin (M-Ex/Im) computes the gap between chosen and rejected responses using either external reward models or implicit DPO rewards. For (2) and (3), we segment the data into P (most positive pairs), Z (close to zero pairs), and N (most negative pairs) subsets according to margin values. Specifically, previous work (Wu et al., 2024b) posits that "hard" preference pairs (where chosen and rejected samples are highly similar) are more beneficial for training, and we use the IFD-Z to quantify this scheme and call it Low-Gap. For the iterative DPO setting, we compare our approach against the standard online iterative DPO baseline established by (Xiong et al., 2024; Dong et al., 2024) and run for three rounds, each using 20k prompts sampled from UltraFeedback. We provide source code of our paper in https://github.com/xiangtanshi/DPO-Data-Selection.
Table 2: GPT-4 judged win rates for Llama-3.2-3B models fine-tuned with DPO on subsets selected by various data selection strategies. For every strategy and benchmark (TL;DR, $H$, UltraFeedback (UF)), 2,000 preference samples were selected. Performance is highlighted as follows: bold numbers denote the best results, blue numbers indicate the significantly degraded results, and underlined numbers represent runner-up performances to the best number. P, Z, and N denote the most positive, near-zero, and most negative selection principles.